A Bibliography of Publications by, and about, Peter Henrici

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

29 November 2023 Version 1.06

Abstract

This bibliography records publications of Peter Henrici.

Title word cross-reference

$1.35 [Hen57a]. $12.50 [Hen57b]. _{2d} [BBH^{+}72]. _{4h} [BBH^{+}72]. ∫ e^{b(x+a cosx)} dx [Hen52c, Hen52c].

11 [Hen64c]. 1867-corrected [Hen57b]. 1967 [DH69]. 1971 [Mor71]. 1975 [Nic75]. 1982 [Hen82e]. 1983 [Bir83].

2 [S.78]. 25 [Hen67, Hen77b].

328 [God65].

50 [Bac77].

60th [Bir83].

82-17 [Hen82c].

Abb [Sch66b]. Abbildung [Hen48]. abgerundeten [Hen48]. abscissa [Hen70c]. accelerating [Sad90]. Accélération [Hen79c]. Accumulation [Hen59b, Hen60a]. Addition [Hen55a]. Advanced [BHJ80, CH57a, CH57b, GH57a, GH57b]. Albert [Hen57a]. Algebra [DH69, Bir80, I.72]. algebraic [Hen64a]. algorithm [GH67, Hen58c, Hen63d, HW65, Hen66a, Hen66b, Hen71a, Hen83c, Tho66]. Algorithmen [Hen89]. Algorithms [Hen89, HG69, Hen70b, Hen83a]. analyse [Hen79c]. Analysis [Bac77, Bel03, Boa87, Bur89, Gol84, Haw78, Hen64b, Hen67, Hen72a, Hen72b, Hen74a, Hen77c, HJ80, Hen82d, Hen86a, Hen88a, Hen91, Mor71, S.87, Tha81, Bir80, BH62, Hen55b, Hen74b, Hen79d, Hen83b, Hen83c, Hen83d, Hen83e, Hen87a, Uhl66, Aus88, Boa76, Boa78, D.66, God65, Ham65, Har65, Hei67, How66, Mac65a, Mac65b, Oyi86, Tha77, Wil65]. Analysis-Cauchy [Boa87]. analytic [Hen57e, Hen66a, Hen83d]. Analytical [Hen77b]. Analytische [Hen77b]. Anwendungen [Hen52d, Hen53b, Hen75, Hen76]. aperture [HTW70]. Application [Hen55b, HW87, TA95]. Applications [Hen75, Mor71, Hen52d, Hen53b, Hen63d, Hen76]. Applied [Bir80, Hen74a, Hen77c, Hen86a, Hen86b, Hen88a, Hen91, Sch66c, Bac77, Aus88, Boa76, Boa78, Boa87, Bur89, Oyi86, S.87, Tha77]. approximate [CH60]. approximation [HG69, Hen84c]. Approximations [FHM67, Hen83d]. Arithmetic [Hen75, GH72, Hen71b, Hen80b]. Art [Hen52a]. Aspects [DH69, I.72]. asymptotic [Hen55d, Hen83d]. Asymptotics [Bur89, Hen77c, Hen91, HH75]. asymptotischen [Hen55d]. Automatic [Hen56a]. axially [Hen57c].

B [Bir80, I.72]. Barycentric [Hen79b]. Bateman [Hen57a]. Bd [Bac77]. behavior [Hen83d]. Bemerkung [Hen52c, Hen55d]. Bennett [Hen57a]. Berechnung [Hen50, Hen79a]. Beresford [Hen80a]. Bergman [Hen52a]. Bergmans [Hen52a]. Bessel [Hen52d, Hen53b, Hen57d]. Besselschen [Hen52d, Hen53b]. best [Hen83d]. Bieberbach [Hen87a]. Bierens [Hen57b]. Birthday [Bir83]. Book [Bir80, Boa76, Boa78, Boa87, Bur89, D.66, God65, Gre64, Gre62, H.65, Ham65, Har65, Haw78, Hei67, Hen55c, Hen59a, Hen79c, Hen80a, Hen86b, How66, I.72, Mac65a, Moo62, O.64, Oyi86, S.78, S.87, Sch66a, Tha77, Tha81, VJ83, Wil64, Wil65, ABS^{+}64, BdHF^{+}64, DJK^{+}82, DSH^{+}85, EHdH^{+}65, EKE^{+}81, HZP^{+}66, HMC^{+}66, HB69, HBA71, HHG^{+}71, Hen78, HES^{+}81, Hen82a, HKC^{+}82, Hen82b, HCG^{+}71, KHCW81b, KHCW81a, KHE^{+}68, LHL^{+}79, LKM^{+}84, MHZ^{+}68, SBM^{+}85, SMO^{+}80, SZR^{+}64, TGH^{+}79, TBH^{+}68, TBS^{+}86, ZFL^{+}66]. Boundary [Hen59a]. Bounds [Hen62a, Hen63a, Hen64c, FHM67, Hen65a, Hen70c]. Branges [Hen87a].

Buchbesprechung [Bac77, Sch66b, Sch66c, Uhl66]. Buchbesprechungen [ABS^{+}64, BdHF^{+}64, DJK^{+}82, DSH^{+}85, EHdH^{+}65, EKE^{+}81, HZP^{+}66, HMC^{+}66, HB69, HBA71, HHG^{+}71, Hen78, HES^{+}81, Hen82a, HKC^{+}82, Hen82b, HCG^{+}71, KHCW81b, KHCW81a, KHE^{+}68, LHL^{+}79, LKM^{+}84, MHZ^{+}68, SBM^{+}85, SMO^{+}80, SZR^{+}64, TGH^{+}79, TBH^{+}68, TBS^{+}86, ZFL^{+}66]. Bunch [Bir80]. Bürmann [Hen64a, Hen83e, Hen84a, HW87]. Bürmannsche [Hen84a].

C [Bir80]. calculation [Hen79a, Hen50]. Calculations [Hen77b]. Calculator [Hen67, Hen77b, Hen82d, Haw78, Tha81, Bir80]. card [Hen50]. Cauchy [Boa87, Bur89, S.87, Hen86a]. Certain [Hen57d, Hen64c, HZ68, Hen55e, Hen63a]. champ [HH75]. chances [HH75]. Chess [HH75]. Circular [GH72, Hen71b, HTW70]. citations [Hen82e]. class [Hen60b]. Clausthal [AC76]. Cliffs [Hen80a]. cm [Hen57a, Hen57b]. Co [Hen57b]. coefficients [Hen57e]. Comment [Hen57a, Hen57b, Hen52c]. commuting [BB93]. Compact [Bir80]. Complete [Hen60b]. Complex [Aus88, Bac77, Boa76, Boa78, Boa87, Bur89, FH60, Hen74a, Hen77c, Hen86a, Hen88a, Hen91, Oyi86, S.87, Tha77, GH67, Hen74b, Hen79d, Hen83b, Hen83c, Hen83d, Hen83e, HJ80]. Computation [Ste74, GH67, Hen55b, TA95]. Computational [Aus88, Bac77, Boa76, Boa78, Boa87, Bur89, Hen67, Hen74a, Hen74b, Hen77c, Hen86a, Hen88a, Hen91, Oyi86, S.87, Tha77, Hen79d, Hen83b, Hen83c, Hen83d, Hen83e, Hen87a, Bir80, Haw78, Tha81]. computations [Hen56a, Hen56b, Hen65a]. computers [Bir80]. Computing [FH60, Hen58b]. concerning [Hen83c]. conducted [DH69]. Conference [Mor71]. Conformal [Boa87, Bur89, Hen74a, Hen86a, Hen88a, S.87, Hen83a]. conforming [Hen48]. conjecture [Hen87a]. conjugates [Hen79b]. conjugation [Hen82f]. Construction [Boa87, Bur89, Hen86a, S.87]. Constructive [DH69, I.72]. continuation [Hen66a]. Continued [Bur89, Hen75, Hen77c, Hen91, GH67, Hen65a]. Convergence [CH60, Hen58b, Hen79c]. convergent [HG69, Hen70b]. corners [Hen48]. corrected [Hen57b]. Corrections [Hen66b, Tho66]. Cyclic [FH60, Hen58b, HZ68].

D [BBH^{+}72, BBH^{+}72, HW65, Hen66b, Tho66]. D. [Hen57b]. data [Hen66b, Tho66]. decomposition [Hen71a]. Définies [Hen57b]. Definite [Hen82c, VJ83]. Dejon [I.72]. Demonstrations [Hen82d]. derivatives [TA95]. descent [Hen83b]. determination [GH72, Hen70b, Hen71b]. developments [Hen83c]. diagonalization [BBH^{+}72]. Difference [Gre64, H.65, Hen63b, Hen77d, Her66, O.64, Wil64, Hen58c, Hen63d, Hen83c, Sch66c]. Differential [Far62, Gre62, Hen57a, Hen59a, Hen59b, Hen62b, Moo62, Sch66a, Sch66b, Hen55c, Hen57e, Hen60b, Hen60c, Hen60a, Hen63c, Hen65b]. digital [Hen65b]. d’Intégrales [Hen57b]. Discrete [Boa87, Bur89, Hen62b, Hen86a, S.87, Hen84c, Sch66b, Far62, Gre62, Moo62, Sch66a]. disk [Hen83d]. domain [Hen57c]. Dongarra [Bir80]. Dover [Hen57a]. Dundee [Mor71].

Dundee/Scotland [Mor71]. durch [Hen48, Hen50].

E2720 [BHJ^{+}79]. E2791 [VPC^{+}79]. E2796 [HDC^{+}80, VPC^{+}79]. E2803 [SBL^{+}79]. E2808 [HH81, SBL^{+}79]. ebenen [Hen48]. Ecken [Hen48]. edition [Hen57b]. Editor [Hen66b, Tho66]. Eduard [WKSH79]. eigenfunction [Hen84c]. eigensystems [TA95]. Eigenvalue [Hen80a, BH62]. Eigenvalues [Hen58b, Hen64c, Nic67, FHM67, Hen50, Hen63a, Hen84c]. eigenvectors [CH60]. Eigenwerte [Hen50]. Einige [Hen75, Hen76]. Elementary [BHJ^{+}79, HDC^{+}80, HH81, SBL^{+}79, VPC^{+}79]. Elemente [Hen72a, Hen72b]. Elementes [Uhl66]. Elements [Hen64b, Hei67, Hen72a, Hen72b, D.66, God65, Ham65, Har65, How66, Mac65a, Mac65b, Wil65]. Elliptic [Nic67, FHM67, Hen55c, Hen57e, Hen60b]. Engineers [HJ80]. Englewood [Hen80a]. Entwicklung [Hen55d]. equation [Hen52d, Hen53b, Hen84c]. Equations [Hen59b, Hen62b, Hen55c, Hen57e, Hen60b, Hen60c, Hen60a, Hen63c, Hen65b, Hen70a, Hen83b, Hen57a, Sch66b, Far62, Gre62, Hen59a, Moo62, Sch66a]. Errata [Hen64c]. Error [Hen59b, Hen63b, Hen65a, Hen77d, Hen55d, Hen60c, Hen60a, Hen63c, Hen65b, HP66, Hen80b, Hen82f, PH66, Gre64, Sch66c, H.65, Her66, O.64, Wil64]. Errors [Hen66c]. erster [Hen52a]. Essentials [Hen82d]. Estimate [HZ68]. estimates [HP66, Hen82f, PH66]. evolution [Hen55d]. Exact [BBH^{+}72]. exercise [HH75]. expansion [HW87]. expansions [Hen55e, Hen84c]. Experimental [Hen59b, Hen60a]. Extension [Ben96, Nic67].

Fast [Hen76, Hen79d, Hen89]. favorite [Hen71c]. features [Hen53b]. Fehlerintegrals [Hen55d]. Fields [Hen62a]. Fig [Sch66c, Uhl66]. Finding [HW65]. finite [Hen55g]. first [Hen52a]. Floating [Ste74, Hen80b]. flow [Hen54a]. formal [Hen83e, Hen84a, Hen88b]. formalen [Hen84a]. Formel [Hen84a]. formula [Hen64a, Hen83e, Hen84a]. formulas [Hen79b]. four [BBH^{+}72]. four-membered [BBH^{+}72]. Fourier [Boa87, Bur89, Hen76, S.87, Hen79d, Hen84c, Hen86a]. Fouriertransformation [Hen76]. Fox [Hen59a, Nic67]. Fraction [Hen75, GH67]. Fractions [Bur89, Hen77c, Hen91, Hen65a, HP66, Hen71a, PH66]. frequencies [HTW70]. Fresnel [Hen79a]. Fresnelschen [Hen79a]. function [Bir80, Hen52a, Hen52d, Hen53b, Hen71a]. Functions [Bur89, Hen57d, Hen77c, Hen86a, Hen89, Hen91, GH67, Hen52d, Hen53a, Hen54b, Hen55a, Hen55c, Hen55e, Hen55f, Hen83d, Boa87, Bur89, S.87]. Fundamental [DH69, I.72]. Funktion [Hen52a]. Funktionen [Hen54b, Hen52d, Hen53b, Hen89]. Funktionentheorie [Hen52d, Hen53b]. Further [Hen52c].

G [Bir80]. geb [Uhl66]. Gebiete [Hen52b]. Gegenbauer [Hen54b, Hen55a]. general [Hen55a, Hen83a]. generalized [Hen57c]. generating [Hen55f]. geodesy [HW87]. German [ABS^{+}64, BdHF^{+}64, DJK^{+}82, DSH^{+}85, EHdH^{+}65, EKE^{+}81, HZP^{+}66, Hen48, Hen50, Hen52a, Hen52d, Hen52b, Hen52c, Hen53b, Hen54b, Hen55d, HMC^{+}66, HB69, HBA71, HHG^{+}71, Hen72a, Hen72b, Hen75, Hen76, Hen77b, HJ80, Hen78, Hen79a, HES^{+}81, Hen82a, HKC^{+}82, Hen82b, Hen82e, Hen84a, Hen89, HCG^{+}71, KHCW81b, KHCW81a, KHE^{+}68, LHL^{+}79, LKM^{+}84, MHZ^{+}68, SBM^{+}85, SMO^{+}80, SZR^{+}64, TGH^{+}79, TBH^{+}68, TBS^{+}86, ZFL^{+}66]. Germany [Nic75]. Gilbert [Hen86b]. guide [Bir80]. Gutknecht [S.78].

Haan [Hen57b]. Hafner [Hen57b]. half [HTW70]. half-space [HTW70]. Hall [Hen80a]. Hans [S.78]. Hans-Rudolf [S.78]. Harry [Hen57a]. Heinz [S.78]. Held [Mor71]. helical [Hen55g]. Henrici [Bac77, Bir80, Boa76, Boa78, Boa87, God65, GV88, Gre64, Gre62, Hei67, S.78, Sch66b, Sch66c, Uhl66, VJ83, Wil64, Wil65, Aus88, Bel03, Ben96, BB93, Bir83, Bur89, D.66, Far62, H.65, Ham65, Har65, Haw78, Hen82e, Her66, How66, I.72, KS90, Mac65a, Mac65b, Moo62, Nic67, O.64, OR08, Oyi86, Rha22, S.87, Sad90, Sch66a, TA95, Tha77, Tha81]. Hermitian [Hen58b]. higher [GH67]. HP [Bir80, Haw78, Tha81, Hen67, Hen77b]. HP-25 [Haw78, Tha81, Hen77b]. hypercube [Hen84c]. Hypergeometric [Hen57d, Hen87b, Hen88b].

I. [Hen57e]. IBM [DH69]. II [Hen72b, Hen83c]. III [Hen83d]. Inc. [Sch66c]. incomplete [Hen71a]. Industrial [Bir80]. Inequality [Hen61a, Hen61b]. infinite [BBH^{+}72]. Ingenieure [HJ80]. Initial [Hen59b, Hen60c, Hen60a]. Integral [Bur89, Hen57d, Hen77c, Hen82c, Hen91, Hen52a, Hen55d, Hen71c, VJ83]. Integrale [Hen52d, Hen53b, Hen79a]. Integraloperator [Hen52a]. Integrals [Boa87, Bur89, Hen86a, S.87, Hen52d, Hen53b, Hen79a]. Integrals-Construction [Boa87]. Integration [Bur89, Hen57a, Hen74a, Hen88a, Hen63c, Hen65b, Hen84b]. international [Nic75]. interpolating [Hen79b]. interpolation [Hen82f]. Interscience [Bur89]. Interval [Hen75, Nic75]. Introduction [Hen86b]. Inverses [Hen62a]. Involving [Hen57d, Hen55e, Hen60c]. Iterates [Hen62a]. IV [Hen83e].

J [Bir80, Hen64c]. Jacobi [CH60, FH60, Hen55e, Hen55f, Hen58b, HZ68]. Jahren [Hen82e]. Jahrfeier [AC76]. John [Bac77, Bir80, Boa87, God65, Sch66b, Sch66c, Uhl66]. joint [BB93]. June [DH69].

Kantorovich [Hen61a, Hen61b]. Karlsruhe [Nic75]. Kernel [Hen55c]. Kettenbruchtheorie [Hen75]. kind [Hen52a]. Kleine [Hen55d]. Kleinere [Hen52b]. Komplexe [HJ80]. konforme [Hen48].

Kreisscheibenarithmetik [Hen75]. Kummer [Hen58a].

L [Hen59a]. Laboratory [DH69]. Lagrange [Hen84a, Hen64a, Hen83e, Hen84a, HW87]. Lagrange-Bürmannsche [Hen84a]. Lauchli [S.78]. Legendre [Hen55a]. Letter [Hen66b, Tho66]. like [HTW70]. linear [Bir80]. LINPACK [Bir80]. Location [Bur89, Hen74a, Hen88a]. Lochkartenmaschinen [Hen50]. London [Bac77, Sch66b, Sch66c, Uhl66]. Lösung [Hen48].

m [Sch66b, Uhl66]. machines [Hen50]. Mapping [Bur89, Hen74a, Hen88a, Hen48, Hen83a]. Maps [Boa87, Bur89, Hen86a, S.87]. Maps-Univalent [Boa87]. March [GV88, Mor71]. Martin [S.78]. Mathematical [Hen52b, Hen61a, Hen55c, HW87]. Mathematics [Bir80, Sch66c, Nic75, Hen86b]. Mathematik [AC76, S.78, Hen52b]. Mathieu [Hen52d, Hen53b]. Mathieuschen [Hen52d, Hen53b]. Matrices [Hen58b, Hen62a, Hen64c, BB93, Hen50, Hen63a]. Matrix [FH60, BBH^{+}72, Ben96, BH62]. Matrizen [Hen50]. May [Nic75]. Mehler [Hen71c]. membered [BBH^{+}72]. Method [FH60, Ben96]. Methoden [AC76]. Methods [Far62, Gre62, Hen58b, Hen62b, Hen63b, Hen70a, Hen77d, Moo62, Sch66a, Sch66b, Bir80, CH60, Hen55b, Hen79d, Hen84c, Hen83b, Gre64, H.65, Her66, O.64, Wil64]. Milne [Hen57a]. minimization [Bir80]. Minor [Hen52b]. model [Hen80b]. Models [Hen66c]. Moderne [AC76]. modified [Rha22]. molecules [BBH^{+}72]. Moler [Bir80, Nic67]. multivariate [HW87]. My [Hen71c].

N [Hen57e, Hen80a]. Nash [Bir80]. net [Uhl66]. Neumann [Hen53a]. NJ [Hen80a]. no [Hen64c]. Non [Hen62a]. Non-Normal [Hen62a]. Normal [Hen62a]. Norms [HZ68]. Note [Hen77a, KS90, Hen55d]. Notes [Hen61a]. Nouvelles [Hen57b]. numbers [Hen56b]. Numerical [D.66, God65, Gol84, Ham65, Har65, Hei67, Hen59a, Hen59b, Hen64b, Hen82d, How66, Mac65a, Mac65b, Mor71, Uhl66, Wil65, Bir80, Hen55b, Hen60c, Hen60a, Hen63c, Hen66b, Hen72a, Hen72b, Hen79a, Hen84b, Tho66, Hen57a]. numérique [Hen79c]. Numerische [S.78]. Numerischen [Hen72b, AC76, Hen72a, Hen79a].

Obituary [GV88]. off [Hen60c]. operator [Hen52a]. Operators [HZ68, Nic67, FHM67]. optimization [Rha22]. order [Hen60c]. Ordinary [Far62, Gre62, Hen59a, Hen59b, Hen62b, Moo62, Sch66a, Sch66b, Hen60c, Hen60a, Hen63c, Hen65b]. osculation [Hen83a].

P [Bac77, Gre62, I.72, Sch66b, Sch66c, Uhl66, Hen57a, Hen57b]. P. [Gre64, Hei67]. Padé [Hen84c]. Paper [Nic67]. Parlett [Hen80a]. partial [Hen57e, Hen60b, Hen71a]. Periodic [Hen89]. periodische [Hen89]. Peter [Aus88, Bir80, Boa76, Boa78, Boa87, Bur89, D.66, Far62, God65, H.65, Ham65, Har65, Haw78, Her66, How66, Mac65a, Mac65b, Moo62, O.64, Oyi86, S.78, S.87, Sch66a, Tha77, Tha81, VJ83, Wil64, Wil65, Bir83, GV88, Hen82e, OR08]. Philadelphia [Bir80]. physics [Hen55c]. planar [Hen48]. plane [GH67]. Pocket [Haw78, Hen67, Hen82d, Tha81, Bir80]. Point [Hen59a, Ste74, Hen55b]. Pointwise [Hen82f]. Poisson [Hen84c]. polynomial [GH72, HW65, HG69, Hen70a, Hen70c, Hen70b, Hen71b, Hen83b]. polynomials [Hen55e, Hen55f, Hen79b]. potential [Hen48]. Potentialproblemen [Hen48]. potentials [Hen57c]. Potenzreihen [Hen84a]. Power [Bur89, Hen74a, Hen88a, Hen56a, Hen83e, Hen84a]. pp [Bir80, God65]. pp. [Bir80, Hen80a]. Preis [Uhl66]. Prentice [Hen80a]. Prentice-Hall [Hen80a]. Presentations [Her66, Mac65b]. Price [Hen57a, Hen57b]. Principal [FH60]. Probabilistic [Hen66c]. Problem [Hen82c, BBH^{+}72, Hen80a]. Problems [BHJ80, Bel03, BHJ^{+}79, CH57a, CH57b, GH57a, GH57b, Hen59a, Hen59b, Hen63c, HDC^{+}80, HH81, SBL^{+}79, VPC^{+}79, BH62, Hen48, Hen60c, Hen60a]. proceedings [DH69, Nic75]. Product [Hen88b, KS90, Hen53a, Hen58a, Hen87b]. proof [Hen64a, Hen71c, Hen87a]. Propagation [Gre64, H.65, Hen63b, Hen66c, Hen77d, Her66, O.64, Sch66c, Wil64, Hen60c, Hen63c, Hen65b, Hen80b]. Publications [Far62, Hen52b, Hen57a, Her66, Mac65b]. Publishing [Hen57b]. puckering [BBH^{+}72]. punch [Hen50].

Q [HW65, Hen66b, Tho66]. Quasicyclic [Hen58b]. quotient [Hen58c, Hen63d, Hen83c]. quotient-difference [Hen83c].

R [Bir80]. radiation [Hen55b]. rational [Hen56b, Hen71a]. Rechenverfahren [Hen77b]. reflected [Hen55b]. regularity [Hen57c]. Reihen [Hen52d, Hen53b]. Remark [Hen84b]. Remarks [Hen61a, Hen61b]. Report [Hei67]. Representation [Hen57d]. Research [DH69]. Review [Bir80, Boa76, Boa78, Boa87, D.66, God65, Gre64, Gre62, H.65, Ham65, Har65, Haw78, Hen55c, Hen59a, Hen79c, Hen80a, Hen86b, How66, I.72, Mac65a, Moo62, O.64, Oyi86, S.78, S.87, Sch66a, Tha77, Tha81, VJ83, Wil64, Wil65]. Reviews [ABS^{+}64, Aus88, BdHF^{+}64, Bur89, DJK^{+}82, DSH^{+}85, EHdH^{+}65, EKE^{+}81, HZP^{+}66, HMC^{+}66, HB69, HBA71, HHG^{+}71, Hen78, HES^{+}81, Hen82a, HKC^{+}82, Hen82b, HCG^{+}71, KHCW81b, KHCW81a, KHE^{+}68, LHL^{+}79, LKM^{+}84, MHZ^{+}68, SBM^{+}85, SMO^{+}80, SZR^{+}64, TGH^{+}79, TBH^{+}68, TBS^{+}86, ZFL^{+}66]. Riemann [Hen52a]. Riemannsche [Hen52a]. ring [BBH^{+}72]. rotation [BBH^{+}72]. rotation-ring-puckering [BBH^{+}72]. rotational [Hen54a]. round [Hen60c]. round-off [Hen60c]. rounded [Hen48]. rounding [Hen80b]. Roundoff [Hen66c]. Rudolf [S.78]. Rüschlikon [DH69]. Rutishauser [S.78].

S [Bac77, Sch66b, Sch66c, Uhl66]. Saff [Hen77a]. Salutation [Bir83]. scharfen [Hen48]. Schnelle [Hen89]. schnellen [Hen76]. Schwarz [S.78]. Scotland [Mor71]. search [Hen70a]. second [Hen60c]. September [GV88, Bir83]. sequences [Ben96, Sad90]. Series [Bur89, Hen57d, Hen74a, Hen88a, Sch66c, Hen52d, Hen53b, Hen53a, Hen55e, Hen56a, Hen58a, Hen83e, Hen84a, Hen87b, Hen88b]. sharp [Hen48]. SIAM [Hen64c, Sch66c]. Sign [BH62]. simultaneous [HG69, Hen70b]. Singular [Bel03, Hen60b]. Sloshing [HTW70]. Small [Hen55d]. Society [Bir80]. Sohn [Uhl66]. Solution [Hen59a, Hen59b, BBH^{+}72, Hen60c, Hen60a]. Solutions [BHJ80, BHJ^{+}79, CH57a, CH57b, GH57a, GH57b, HDC^{+}80, HH81, SBL^{+}79, VPC^{+}79, Hen60b]. solving [Hen48, Hen70a]. Some [Hen63d, Hen75, Hen76]. Sons [Bir80, God65, Sch66c, Bac77, Sch66b]. source [Hen55b]. space [HTW70]. Special [Bur89, Hen77c, Hen91, Hen52d, Hen53b]. spectra [BB93]. Spectral [Hen62a]. Speed [Hen58b]. spezielle [Hen52d, Hen53b]. springs [Hen55g]. stability [Hen63c, Hen70c]. stable [Hen70c]. Stewart [Bir80]. Stiefel [WKSH79]. Stieltjes [HP66, PH66]. Strang [Hen86b]. strip [HTW70]. strip-like [HTW70]. Studies [Gol84, Hen59b, Hen60a]. subroutine [Hen56b]. survey [Hen57e]. Switzerland [DH69]. Sydney [Bac77, Uhl66]. Symmetric [Hen80a, Hen57c]. symposium [DH69, Nic75]. Systems [Hen59b, Hen60b, Hen60a, Hen83e].

Tab [Sch66b]. Tables [Hen57b]. Tagung [AC76]. Taschenrechner [Hen77b]. Technische [AC76]. Test [Hen66c]. their [Hen79b]. Theorem [DH69, I.72, KS90, BB93, Hen77a, Hen87b]. theorems [Hen55a, Hen88b]. Theoretical [Hen59b, Hen60a]. Theory [Hen75, Hen52d, Hen53b, Hen57e, Hen83a]. thickness [Hen55g]. Three [Boa87]. Topics [Hen83b, Hen83c, Hen83d, Hen83e]. Toronto [Bac77]. transcendental [GH67]. Transform [Hen76]. Transformation [Bel03, Rha22, Sad90, TA95]. Transforms [Bur89, Hen77c, Hen91]. Tridiagonal [Hen64c, Hen63a]. trigonometric [Hen79b, Hen82f]. Triple [KS90, Hen87b]. Truncation [HP66, PH66]. tube [Hen54a]. Two [Hen59a, Hen61a, Hen61b, Hen53a, Hen55b, Hen58a]. Two-Point [Hen59a].

u [Sch66b, Sch66c]. Uniformly [HG69, Hen70b]. unit [Hen83d]. Univalent [Boa87, Bur89, Hen86a, S.87]. Universität [AC76]. Upper [Hen70c]. User [Bir80].

Value [Hen59b, Hen60c, Hen60a]. Values [FH60, Hen62a]. Varga [Hen77a]. Variable [Far62, Gre62, Hen62b, Moo62, Sch66a, Sch66b]. Variation [Hen62a]. vector [Sad90]. Vekua [Hen57e]. Veröffentlichungen [Hen52b]. version [Rha22]. view [Hen87a]. viscous [Hen54a]. Vol [Bur89, Oyi86, S.87]. Vols [S.78]. Volume [Boa87, Hen74a, Hen77c, Hen86a, Hen88a, Hen91]. Vorlesungen [S.78].

W [Bir80]. wave [BH62, Hen52d, Hen53b]. Weitere [Hen52c]. Wellengleichung [Hen52d, Hen53b]. West [Nic75]. Whittaker [Hen52d, Hen53b, Hen53a, Hen55e]. Whittakerschen [Hen52d, Hen53b]. Wiley [Bac77, Bir80, Boa87, Bur89, God65, Sch66b, Sch66c, Uhl66]. Wiley-Interscience [Bur89]. William [Hen57a].

x [Bir80]. xi [Sch66b]. xiv [Hen57b]. xv [Uhl66].

years [Hen82e]. York [Bac77, Bir80, God65, Hen57a, Hen57b, Sch66b, Sch66c, Uhl66]. York/London [Sch66b, Sch66c]. York/London/Sydney [Uhl66].

Zeros [Bur89, Hen74a, Hen88a, GH72, HW65, HG69, Hen70b, Hen71b]. Zitate [Hen82e]. zur [Hen55d, Hen50, Hen52d, Hen53b, Hen79a]. Zürich [DH69]. Zürich-Rüschlikon [DH69].

References

Alder:1964:BGB

[ABS^{+}64] K. Alder, E. Baldinger, W. Saxer, P. Henrici, B. Federer, H. Ziegler, W. Czaja, and E. Roth. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 15:93– 96, January 1964. CODEN ZAMPDB. ISSN 0044-2275 (print),

           1420-9039 (electronic). URL http://adsabs.harvard.edu/abs/
           1964ZaMP...15...93A.

Albrecht:1976:MMN

[AC76] J. Albrecht and L. Collatz, editors. Moderne Methoden der numerischen Mathematik (Tagung 200–Jahrfeier, Technische Universität Clausthal, Clausthal, 1975), volume 32 of International Series in Numerical Mathematics. Birkhäuser Verlag, Basel, Switzerland, 1976. ISBN 3-0348-5501-X. URL https://link.

           springer.com/book/10.1007/978-3-0348-5501-3.

Auslander:1988:RAC

[Aus88] Louis Auslander. Reviews: Applied and Computational Complex Analysis, by Peter Henrici. American Mathematical Monthly, 95 (4):366, April 1988. CODEN AMMYAE. ISSN 0002-9890 (print),

           1930-0972 (electronic). URL https://www.jstor.org/stable/
           2323585.

Bachmann:1977:BHP

[Bac77] K.-H. Bachmann. Buchbesprechung: Henrici, P., Applied and Computational Complex Analysis, Bd. I, 682 S., New York– London–Sydney–Toronto. John Wiley & sons. 1974. £13,50. Zeitschrift für Angewandte Mathematik und Mechanik, 57(6):352, 1977. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Bhatia:1993:HTJ

[BB93] Rajendra Bhatia and Tirthankar Bhattacharyya. A Henrici theorem for joint spectra of commuting matrices. Proceedings of the American Mathematical Society, 118(1):5–14, May 1993. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL

https://www.jstor.org/stable/2160002.

Baltagi:1972:ESR

[BBH^{+}72] F. Baltagi, A. Bauder, P. Henrici, T. Ueda, and H. H. Günthard. Exact solution to the rotation-ring-puckering problem of D_{4h}–D_{2d} four-membered ring molecules by infinite matrix diagonalization. Molecular Physics, 24:945–968, 1972. CODEN MOPHAM. ISSN 0026-8976 (print), 1362-3028 (electronic). URL http://adsabs.

harvard.edu/abs/1972MolPh..24..945B.

Blanc:1964:B

[BdHF^{+}64] C. Blanc, J. P. den Hartog, M. Fierz, P. Henrici, J. Ackeret, E. Baldinger, H. Badr, E. Bas, P. Profos, W. Saxer, E. Stiefel, T. M. Gallie, H. Ziegler, and W. Gerber. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 15: 567–572, September 1964. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://adsabs.harvard.

edu/abs/1964ZaMP...15..567B.

Bellalij:2003:AHT

[Bel03] Mohammed Bellalij. Analysis of Henrici’s transformation for singular problems. Numerical Algorithms, 33(1–4):65–82, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/46/6/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/46/6/
fulltext.pdf.

Benchiboun:1996:EHM

[Ben96] M. D. Benchiboun. Extension of Henrici’s method to matrix sequences. Journal of Computational and Applied Mathematics, 75(1):1–21, November 12, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0377042796000313.

Brown:1962:SWA

[BH62] K. M. Brown and P. Henrici. Sign wave analysis in matrix eigenvalue problems. Mathematics of Computation, 16:291–300, 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boas:1979:PSS

[BHJ^{+}79] Ralph P. Boas, P. Henrici, A. A. Jagers, Leon Gerber, Stephen V. Noltie, H. Kestelman, Otto G. Ruehr, Michael Vowe, Robert E. Shafer, and C. C. Rousseau. Problems and solutions: Solutions of elementary problems: E2720. American Mathematical Monthly, 86 (8):706–707, 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Baillie:1980:PSS

[BHJ80] Robert Baillie, P. Henrici, and Richard Johnsonbaugh. Problems and solutions: Solutions of advanced problems: 6241. American Mathematical Monthly, 87(6):496–498, 1980. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://links.jstor.org/sici?sici=0002-9890(198006/
07)87:6<496:6>2.0.CO;2-F&origin=MSN.

Birkhoff:1980:BRC

[Bir80] Garrett Birkhoff. Book review: Computational analysis with the HP 25 pocket calculator, by Peter Henrici, John Wiley, New York, 1977, 280 pp.; Compact numerical methods for computers: linear algebra and function minimization, by J. C. Nash, John Wiley & Sons, New York, 1979, x + 227 pp.; LINPACK: User’s guide, by J. J. Dongarra, J. R. Bunch, C. B. Moler, and G. W. Stewart, Society for Industrial and Applied Mathematics, Philadelphia, 1979 368 pp. Bulletin of the American Mathematical Society, 2(3):503–505, May 1980. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL https:/

/www.ams.org/journals/bull/1980-02-03/S0273-0979-1980-
14785-0/S0273-0979-1980-14785-0.pdf.

Birkhoff:1983:SPH

[Bir83] Garrett Birkhoff. Salutation to Peter Henrici on his 60th birthday: September 13, 1983. SIAM Journal on Numerical Analysis, 20(6): ??, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print),

1095-7170 (electronic). URL https://www.jstor.org/stable/

pdf/2157141.pdf. Volume 20(6) is dedicated to Peter Henrici.

Boas:1976:BRB

[Boa76] R. P. Boas. Book review: Applied and Computational Complex Analysis (Peter Henrici). SIAM Review, 18(3):506–510, July 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200

(electronic). URL https://www.jstor.org/stable/2028655;
https://www.jstor.org/stable/pdf/10.2307/2028655.pdf.

Boas:1978:BRB

[Boa78] R. P. Boas. Book review: Applied and Computational Complex Analysis (Peter Henrici). SIAM Review, 20(4):871, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL https://www.jstor.org/stable/2029767.

Boas:1987:BRB

[Boa87] R. P. Boas. Book review: Applied and Computational Complex Analysis. Volume Three: Discrete Fourier Analysis-Cauchy Integrals-Construction of Conformal Maps-Univalent Functions (Peter Henrici. John Wiley). SIAM Review, 29(1):138, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL https://www.jstor.org/stable/pdf/2030950.
pdf.

Burckel:1989:BRA

[Bur89] R. B. Burckel. Book reviews: Applied and Computational Complex Analysis. Wiley-Interscience. Vol. 1: Power Series, Integration, Conformal Mapping, Location of Zeros. Vol. 2: Special Functions, Integral Transforms, Asymptotics, Continued Fractions. Vol. 3: Discrete Fourier Analysis, Cauchy Integrals, Construction of Conformal Maps, Univalent Functions by Peter Henrici. American Scientist, 77(4):404, 1989. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.org/

stable/27855908.

Carlitz:1957:APSa

[CH57a] Leonard Carlitz and Peter Henrici. Advanced problems and solutions: Solutions: 4686. American Mathematical Monthly, 64(4):

280, April 1957. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL https://www.jstor.org/stable/
2310448.

Carlitz:1957:APSb

[CH57b] Leonard Carlitz and Peter Henrici. Advanced problems and solutions: Solutions: 4686. American Mathematical Monthly, 64(4): 280, 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Causey:1960:CAE

[CH60] Robert L. Causey and Peter Henrici. Convergence of approximate eigenvectors in Jacobi methods. Numerische Mathematik, 2 (1):67–78, December 1960. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/BF01386211.

D:1966:BRB

[D.66] H. A. D. Book review: Elements of Numerical Analysis, by Peter Henrici. Journal of the American Statistical Association, 61 (313):295–296, March 1966. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283084.

Dejon:1969:CAF

[DH69] Bruno Dejon and Peter Henrici, editors. Constructive Aspects of the Fundamental Theorem of Algebra: proceedings of a symposium conducted at the IBM Research Laboratory, Zürich-Rüschlikon, Switzerland, June 5–7, 1967. Wiley, New York, NY, USA, 1969. ISBN 0-471-20300-9. LCCN QA212 .C65.

Doring:1982:BGB

[DJK^{+}82] B. Döring, H. Jarchow, U. Kirchgraber, A. Huber, G. Schüler, P. Henrici, M. Gutknecht, W. Gander, U. Krengel, G. Mazziotto, H. Läuchli, F. Troyon, H. Melchior, A. Buck, P. Minkowski, and C. Cercignani. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 33:835–842, November 1982. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1982ZaMP...
33..835D.

Droz:1985:BGB

[DSH^{+}85] M. Droz, R. Stierlin, J. Hersch, M. Gutknecht, P. Henrici, R. Boutellier, J. J. Burckhardt, and W. Gander. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 36:784–788, September 1985. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://adsabs.harvard.

edu/abs/1985ZaMP...36..784D.

Epprecht:1965:BGB

[EHdH^{+}65] W. Epprecht, P. Henrici, P. de Haller, E. Baldinger, I. Gränacher, H. Weber, W. Saxer, P. Schmid, and W. Schumann. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 16:309–312, March 1965. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://adsabs.

harvard.edu/abs/1965ZaMP...16..309E.

Epprecht:1981:BGB

[EKE^{+}81] W. Epprecht, U. Kirchgraber, E. Engeler, R. Jost, P. Henrici, E. Bas, G. Schüler, J. Marti, K. Hazeghi, R. D. Neuchâtel, and H. Brunner. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 32:768–772, November 1981. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1981ZaMP...
32..768E.

Farnell:1962:RPD

[Far62] A. B. Farnell. Recent publications: Discrete Variable Methods in Ordinary Differential Equations, by Peter Henrici. American Mathematical Monthly, 69(9):935–936, November 1962. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

https://www.jstor.org/stable/2311275.

Forsythe:1960:CJM

[FH60] G. E. Forsythe and P. Henrici. The cyclic Jacobi method for computing the principal values of a complex matrix. Transactions of the American Mathematical Society, 94:1–23, 1960. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Fox:1967:ABE

[FHM67] L. Fox, P. Henrici, and C. Moler. Approximations and bounds for eigenvalues of elliptic operators. SIAM Journal on Numerical Analysis, 4:89–102, 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gallagher:1957:APS

[GH57a] Patrick Gallagher and Peter Henrici. Advanced problems and solutions: Solutions: 4717. American Mathematical Monthly, 64 (9):682, November 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/2309985.

Golomb:1957:APS

[GH57b] S. W. Golomb and Peter Henrici. Advanced problems and solutions: Solutions: 4720. American Mathematical Monthly, 64(10): 749–750, December 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/2309771.

Gargantini:1967:CFA

[GH67] I. Gargantini and P. Henrici. A continued fraction algorithm for the computation of higher transcendental functions in the complex plane. Mathematics of Computation, 21:18–29, 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gargantini:1971:CAD

[GH72] Irene Gargantini and Peter Henrici. Circular arithmetic and the determination of polynomial zeros. Numerische Mathematik, 18(4):305–320, 1971/72. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/BF01404681.

Godfrey:1965:BRE

[God65] M. D. Godfrey. Book review: Elements of Numerical Analysis, by Peter Henrici (John Wiley and Sons, New York, 1964), pp. 328. Naval Research Logistics Quarterly, 12(2):195, June 1965. CODEN NRLQAR. ISSN 0028-1441 (print), 1931-9193 (electronic).

Golub:1984:SNA

[Gol84] Gene H. Golub, editor. Studies in Numerical Analysis, volume 24 of Studies in mathematics. Mathematical Association of America, Washington, DC, USA, 1984. ISBN 0-88385-126-1 (v. 1), 0-88385-100-8 (set). x + 415 pp. LCCN QA297 .S83 1984.

Greenspan:1962:BRB

[Gre62] Donald Greenspan. Book review: Discrete Variable Methods in Ordinary Differential Equations (P. Henrici). SIAM Review, 4(3): 261–262, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenberg:1964:BRP

[Gre64] H. J. Greenberg. Book review: P. Henrici, Error Propagation for Difference Methods. Journal of Applied Mechanics, 31(4):735, December 1964. CODEN JAMCAV. ISSN 0021-8936 (print),

1528-9036 (electronic). URL http://adsabs.harvard.edu/abs/
1964JAM....31..735H.

Golub:1988:OPH

[GV88] Gene H. Golub and Richard S. Varga. Obituary to Peter Henrici (13. September 1923–13. March 1987). Numerische Mathematik, 52(5):481–482, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

H:1965:EPD

[H.65] A. S. H. Book review: Error Propagation for Difference Methods, by Peter Henrici. Mathematics of Computation, 19(89):152, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL https://www.jstor.org/stable/2004119.

Hamming:1965:ENA

[Ham65] R. W. Hamming. Book review: Elements of Numerical Analysis, by Peter Henrici. Mathematics of Computation, 19(91):516–517, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL https://www.jstor.org/stable/2003710.

Harrison:1965:BRBb

[Har65] T. K. Harrison. Book review: Elements of Numerical Analysis, by Peter Henrici. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):252–253, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2986934.

Hawthorn:1978:BRC

[Haw78] John Hawthorn. Book review: Computational Analysis with the HP-25 Pocket Calculator, by Peter Henrici. Mathematical Gazette, 62(420):142–143, June 1978. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.jstor.org/

stable/pdf/3617691.pdf.

Henrici:1969:BGB

[HB69] P. Henrici and H. Bühlmann. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 20:280, March 1969. CODEN ZAMPDB. ISSN 0044-2275 (print),

1420-9039 (electronic). URL http://adsabs.harvard.edu/abs/
1969ZaMP...20..280H.

Henrici:1971:BGBa

[HBA71] P. Henrici, C. Blatter, and K. Alder. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 22:215– 216, January 1971. CODEN ZAMPDB. ISSN 0044-2275 (print),

1420-9039 (electronic). URL http://adsabs.harvard.edu/abs/
1971ZaMP...22..215H.

Hersch:1971:BGB

[HCG^{+}71] J. Hersch, H. Carnal, M. Gutknecht, C. Blatter, and P. Henrici. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 22:807–808, July 1971. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://

adsabs.harvard.edu/abs/1971ZaMP...22..807H.

Henrici:1980:PSS

[HDC^{+}80] P. Henrici, Charles A. DeCarlucii, Jeffrey Mitchell Cohen, Michael J. Dixon, S. J. Goodenough, T. M. Mills, A. A. Jagers, L. Kuipers, L. Van Hamme, and Pei Yuan Wu. Problems and solutions: Solutions of elementary problems: E2796. American Mathematical Monthly, 87(9):824, 1980. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://links.jstor.org/sici?sici=0002-9890(198012)
87:10<824:E>2.0.CO;2-O&origin=MSN.

Heinrich:1967:BRP

[Hei67] H. Heinrich. Book report: P. Henrici, Elements of Numerical Analysis. Computing, 2(2):171, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Henrici:1948:LEP

[Hen48] Peter Henrici. Über die Lösung von ebenen Potentialproblemen mit scharfen und abgerundeten Ecken durch konforme Abbildung. (German) [On solving planar potential problems with sharp and rounded corners by conforming mapping]. Diplomarbeit, ETH, Zürich, Switzerland, 1948. 161 pp.

Henrici:1950:BEM

[Hen50] Peter Henrici. Zur Berechnung der Eigenwerte von Matrizen durch Lochkartenmaschinen. (German) [Calculation of the eigenvalues of matrices by punch card machines]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 1:185–189, 1950. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (electronic).

Henrici:1952:BIE

[Hen52a] Peter Henrici. Bergmans Integraloperator erster Art und Riemannsche Funktion. (German) [Bergman’s integral operator of the first kind and Riemann function]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 3(3):228–232, May 1952. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (electronic). URL https://link.

springer.com/article/10.1007/BF02008829.

Henrici:1952:KVG

[Hen52b] Peter Henrici. Kleinere Veröffentlichungen aus dem Gebiete der Mathematik. (German) [Minor mathematical publications]. Report, ETH, Zürich, Switzerland, 1952.

Henrici:1952:WBG

[Hen52c] Peter Henrici. Weitere Bemerkung zu ∫ e^{b(x+a cos x)} dx. (German) [Further comment on ∫ e^{b(x+a cosx)} dx]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 3(6):466–468, November 1952. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (electronic). URL https://link.

springer.com/article/10.1007/BF02025574.

Henrici:1952:FWA

[Hen52d] Peter Henrici. Zur Funktionentheorie der Wellengleichung: mit Anwendungen auf spezielle Reihen und Integrale mit Besselschen, Whittakerschen und Mathieuschen Funktionen. (German) [On function theory of the wave equation: with applications to special series and integrals with Bessel, Whittaker and Mathieu functions]. Dr. sc. math., ETH, Zürich, Switzerland, 1952.

Henrici:1953:NSP

[Hen53a] Peter Henrici. A Neumann series for the product of two Whittaker functions. Proceedings of the American Mathematical Society, 4(2): 329–334, April 1953. CODEN PAMYAR. ISSN 0002-9939 (print),

1088-6826 (electronic). URL https://www.jstor.org/stable/
pdf/10.2307/2031812.pdf.

Henrici:1953:FWA

[Hen53b] Peter Henrici. Zur Funktionentheorie der Wellengleichung. Mit Anwendungen auf spezielle Reihen und Integrale mit Besselschen, Whittakerschen und Mathieuschen Funktionen. (German) [On function theory of the wave equation. With applications to special series and integrals with Bessel, Whittaker and Mathieu features]. Commentarii Mathematici Helvetici (Zurich), 27(1):235–293, December 1953. CODEN COMHAX. ISSN 0010-2571 (print), 1420-

8946 (electronic). URL https://link.springer.com/article/
10.1007/BF02564564.

Henrici:1954:RVF

[Hen54a] Peter Henrici. On rotational viscous flow through a tube. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 5(6):511–516, November 1954. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01601218.

Henrici:1954:FGG

[Hen54b] Peter Henrici. Über die Funktionen von Gegenbauer. (German) [On Gegenbauer functions]. Archiv der Mathematik, 5(1–3):92–98, April 1954. CODEN ACVMAL. ISSN 0003-889X (print), 1420-

8938 (electronic). URL https://link.springer.com/article/

10.1007%2FBF01899322. Herrn Professor Dr. Alexander Ostrowski zum 60. Geburtstag gewidmet.

Henrici:1955:ATG

[Hen55a] Peter Henrici. Addition theorems for general Legendre and Gegenbauer functions. Archive for Rational Mechanics and Analysis, 4: 983–1018, 1955. CODEN AVRMAW. ISSN 0003-9527 (print), 1432-0673 (electronic).

Henrici:1955:ATM

[Hen55b] Peter Henrici. Application of two methods of numerical analysis to the computation of the reflected radiation of a point source. Journal of the Washington Academy of Sciences, 45:38–45, 1955. CODEN JWASA3. ISSN 0043-0439.

Henrici:1955:BRK

[Hen55c] Peter Henrici. Book review: Kernel functions and elliptic differential equations in mathematical physics. Bulletin of the American Mathematical Society, 61(6):596–600, 1955. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Henrici:1955:KBA

[Hen55d] Peter Henrici. Kleine Bemerkung zur asymptotischen Entwicklung des Fehlerintegrals. (German) [Small note on the asymptotic evolution of the error integral]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 6(2):145–146, March 1955. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (electronic). URL https://link.springer.

com/article/10.1007/BF01607304.

Henrici:1955:CSE

[Hen55e] Peter Henrici. On certain series expansions involving Whittaker functions and Jacobi polynomials. Pacific Journal of Mathematics, 5:725–744, 1955. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http://projecteuclid.

org/euclid.pjm/1171984832.

Henrici:1955:GFJ

[Hen55f] Peter Henrici. On generating functions of the Jacobi polynomials. Pacific Journal of Mathematics, 5:923–931, 1955. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL

http://projecteuclid.org/euclid.pjm/1172000956.

Henrici:1955:HSF

[Hen55g] Peter Henrici. On helical springs of finite thickness. Quarterly of Applied Mathematics, 13:106–110, 1955. CODEN QAMAAY. ISSN 0033-569x (print), 1552-4485 (electronic).

Henrici:1956:ACP

[Hen56a] Peter Henrici. Automatic computations with power series. Journal of the ACM, 3(1):10–15, 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735x (electronic).

Henrici:1956:SCR

[Hen56b] Peter Henrici. A subroutine for computations with rational numbers. Journal of the ACM, 3:6–9, 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735x (electronic).

Henrici:1957:CAB

[Hen57a] Peter Henrici. Comment: Albert A. Bennett, William E. Milne, and Harry Bateman, Numerical Integration of Differential Equations, Dover Publications, Inc., New York, 1956, 108 p., 20 cm. Price $1.35. Mathematical Tables and Other Aids to Computation, 11(58):117–118, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.jstor.

org/stable/2002157.

Henrici:1957:CDB

[Hen57b] Peter Henrici. Comment: D. Bierens de Haan, Nouvelles Tables d’Intégrales Définies, Hafner Publishing Co., New York, 1957 (edition of 1867-corrected), xiv + 716 p., 24 cm. Price $12.50. Mathematical Tables and Other Aids to Computation, 11(58):111, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853

(electronic). URL http://www.jstor.org/stable/2002157.

Henrici:1957:DRG

[Hen57c] Peter Henrici. On the domain of regularity of generalized axially symmetric potentials. Proceedings of the American Mathematical Society, 8(1):29–31, February 1957. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL https://

www.jstor.org/stable/pdf/10.2307/2032805.pdf.

Henrici:1957:RCI

[Hen57d] Peter Henrici. On the representation of a certain integral involving Bessel functions by hypergeometric series. Journal of Mathematics and Physics (MIT), 36(1–4):151–156, April 1957. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1957361151.

Henrici:1957:SVT

[Hen57e] Peter Henrici. A survey of I. N. Vekua’s theory of elliptic partial differential equations with analytic coefficients. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 8(3):169–203, May 1957. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (electronic). URL https://

link.springer.com/article/10.1007/BF01600500.

Henrici:1958:PTK

[Hen58a] Peter Henrici. On the product of two Kummer series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10:463–467, 1958. CODEN CJMAAB. ISSN 0008-414X.

Henrici:1958:SCC

[Hen58b] Peter Henrici. On the speed of convergence of cyclic and quasicyclic Jacobi methods for computing the eigenvalues of Hermitian matrices. Journal of the Society for Industrial and Applied Mathematics, 6(2):144–162, June 1958. CODEN JSIMAV.

ISSN 0368-4245. URL https://www.jstor.org/stable/pdf/10.
2307/2098731.pdf.

Henrici:1958:QDA

[Hen58c] Peter Henrici. The quotient–difference algorithm. Applied Mathematics Series (Washington, DC), 49:23–46, 1958. ISSN 1049-4685.

Henrici:1959:BRN

[Hen59a] Peter Henrici. Book review: The Numerical Solution of Two-Point Boundary Problems in Ordinary Differential Equations, by L. Fox. Mathematical Tables and Other Aids to Computation, 13(68):322– 323, October 1959. CODEN MTTCAS. ISSN 0891-6837. URL

https://www.jstor.org/stable/2002815.

Henrici:1959:TES

[Hen59b] Peter Henrici. Theoretical and experimental studies on the accumulation of error in the numerical solution of initial value problems for systems of ordinary differential equations. In ????, editor, Proceedings of the International Conference on Information Processing, UNESCO, pages 36–43. UNESCO, Paris, France, 1959. LCCN ????

Henrici:1960:TES

[Hen60a] P. Henrici. Theoretical and experimental studies on the accumulation of error in the numerical solution of initial value problems for systems of ordinary differential equations. In Information processing, pages 36–44. UNESCO, R. Oldenbourg, and Butterworths, Paris, France, Munich, West Germany, and London, UK, 1960.

Henrici:1960:CSS

[Hen60b] Peter Henrici. Complete systems of solutions for a class of singular elliptic partial differential equations. In Boundary problems in differential equations, pages 19–34. University of Wisconsin Press, Madison, WI, USA, 1960.

Henrici:1960:PRE

[Hen60c] Peter Henrici. The propagation of round-off error in the numerical solution of initial value problems involving ordinary differential equations of the second order. In Symposium on the numerical treatment of ordinary differential equations, integral and integrodifferential equations (Rome, 1960), pages 275–291. Birkhäuser Verlag, Basel, Switzerland, 1960.

Henrici:1961:MNT

[Hen61a] Peter Henrici. Mathematical notes: Two remarks on the Kantorovich inequality. American Mathematical Monthly, 68(9):904– 906, November 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Henrici:1961:TRK

[Hen61b] Peter Henrici. Two remarks on the Kantorovich inequality. American Mathematical Monthly, 68(9):904–906, November 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/pdf/10.2307/
2311698.pdf.

Henrici:1962:BII

[Hen62a] Peter Henrici. Bounds for iterates, inverses, spectral variation, and fields of values of non-normal matrices. Numerische Mathematik, 4(1):24–40, December 1962. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/BF01386294. The paper was presented (in

part) at the [first] Invitational Symposium on Matrix Computation, Gatlinburg, Tennessee, April 23–29, 1961.

Henrici:1962:DVM

[Hen62b] Peter Henrici. Discrete Variable Methods in Ordinary Differential Equations. Wiley, New York, NY, USA, 1962. xi + 407 pp.

Henrici:1963:BEC

[Hen63a] Peter Henrici. Bounds for eigenvalues of certain tridiagonal matrices. Journal of the Society for Industrial and Applied Mathematics, 11(2):281–290, June 1963. CODEN JSIMAV.

ISSN 0368-4245. URL https://www.jstor.org/stable/pdf/10.
2307/2098932.pdf.

Henrici:1963:EPD

[Hen63b] Peter Henrici. Error Propagation for Difference Methods. The SIAM series in applied mathematics. Wiley, New York, NY, USA, 1963. vii + 73 pp. LCCN QA431 .H44. Companion to the author’s Discrete variable methods in ordinary differential equations.

Henrici:1963:PSE

[Hen63c] Peter Henrici. Problems of stability and error propagation in the numerical integration of ordinary differential equations. In Proceedings of the International Congress of Mathematicians (Stockholm 1962), pages 102–113. Institut Mittag-Leffler, Djursholm, Sweden, 1963.

Henrici:1963:SAQ

[Hen63d] Peter Henrici. Some applications of the quotient–difference algorithm. In Proceedings of the Symposium on Applied Mathematics, volume XVI, pages 159–183. American Mathematical Society, Providence, RI, USA, 1963.

Henrici:1964:APL

[Hen64a] Peter Henrici. An algebraic proof of the Lagrange–Bürmann formula. Journal of Mathematical Analysis and Applications, 8:218– 224, 1964. CODEN JMANAK. ISSN 0022-247x.

Henrici:1964:ENA

[Hen64b] Peter Henrici. Elements of Numerical Analysis. Wiley, New York, NY, USA, 1964. xv + 328 pp. LCCN QA297 .H4. Based on the author’s lecture notes for a course entitled “Numerical Mathematical Analysis” at the University of California, Los Angeles.

Henrici:1964:EBE

[Hen64c] Peter Henrici. Errata: “Bounds for Eigenvalues of Certain Tridiagonal Matrices” [J. SIAM 11 (1963), no. 2, 281–290]. Journal of the Society for Industrial and Applied Mathematics, 12(2):497, June 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-

712X (electronic). URL https://www.jstor.org/stable/pdf/
10.2307/2946386.pdf. See [Hen63a].

Henrici:1965:EBC

[Hen65a] Peter Henrici. Error bounds for computations with continued fractions. In Error in Digital Computation, Vol. 2 (Proceedings of the Symposium at the Mathematics Research Center, U.S. Army, University of Wisconsin, Madison, Wisconsin, 1965), pages 39–53. Wiley, New York, NY, USA, 1965.

Henrici:1965:PED

[Hen65b] Peter Henrici. The propagation of error in the digital integration of ordinary differential equations. In Error in Digital Computation, Vol. 1 (Proceedings of the Advanced Seminar Conducted by the Mathematics Research Center, U.S. Army, University of Wisconsin, Madison, Wis, 1964), pages 185–205. Wiley, New York, NY, USA, 1965.

Henrici:1966:AAC

[Hen66a] Peter Henrici. An algorithm for analytic continuation. SIAM Journal on Numerical Analysis, 3(1):67–78, March 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL

https://www.jstor.org/stable/pdf/10.2307/2949686.pdf.

Henrici:1966:LEC

[Hen66b] Peter Henrici. Letter to the Editor: Corrections to numerical data on Q–D algorithm. Communications of the ACM, 9(5):323, May 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [HW65, Tho66].

Henrici:1966:TPM

[Hen66c] Peter Henrici. Test of probabilistic models for the propagation of roundoff errors. Communications of the ACM, 9(6):409–410, June 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Henrici:1967:CAH

[Hen67] Peter Henrici. Computational Analysis with the HP 25 Pocket Calculator. Wiley, New York, NY, USA, 1967. ISBN 0-471-02938-6. xi + 280 pp. LCCN QA297 .H399.

Henrici:1970:MSS

[Hen70a] Peter Henrici. Methods of search for solving polynomial equations. Journal of the ACM, 17(2):273–283, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735x (electronic).

Henrici:1970:UCA

[Hen70b] Peter Henrici. Uniformly convergent algorithms for the simultaneous determination of all zeros of a polynomial. In Studies in Numerical Analysis, 2: Numerical Solutions of Nonlinear Problems (Symposia, SIAM, Philadelphia, PA, 1968), pages 1–8. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1970.

Henrici:1970:UBA

[Hen70c] Peter Henrici. Upper bounds for the abscissa of stability of a stable polynomial. SIAM Journal on Numerical Analysis, 7:538– 544, 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-

7170 (electronic). URL https://www.jstor.org/stable/pdf/
10.2307/2949375.pdf.

Henrici:1971:AID

[Hen71a] Peter Henrici. An algorithm for the incomplete decomposition of a rational function into partial fractions. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 22(4):751–755, July 1971. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (electronic).

Henrici:1971:CAD

[Hen71b] Peter Henrici. Circular arithmetic and the determination of polynomial zeros. In Morris [Mor71], pages 86–92. CODEN LN-MAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 228. URL http://link.springer.com/book/10.1007/
BFb0069442;   http://www.springerlink.com/content/978-
3-540-36976-9.

Henrici:1971:MFP

[Hen71c] Peter Henrici. My favorite proof of Mehler’s integral. American Mathematical Monthly, 78(2):183–185, February 1971. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

https://www.jstor.org/stable/pdf/10.2307/2317631.pdf.

Henrici:1972:ENAa

[Hen72a] Peter Henrici. Elemente der numerischen Analysis. I. (German) [Elements of numerical analysis. I]. Bibliographisches Institut, Mannheim, Germany; Vienna, Austria; Zurich, Switzerland, 1972. 1–228 pp. Übersetzung aus dem Englischen: Götz Alefeld, B·I·- Hochschultaschenbücher, No. 551.

Henrici:1972:ENAb

[Hen72b] Peter Henrici. Elemente der Numerischen Analysis. II. (German) [Elements of numerical analysis. II]. Bibliographisches Institut, Mannheim, Germany; Vienna, Austria; Zurich, Switzerland, 1972. iv + 229–425 pp. Übersetzung aus dem Englischen: Götz Alefeld, B·I·-Hochschultaschenbücher, No. 562.

Henrici:1974:ACC

[Hen74a] Peter Henrici. Applied and Computational Complex Analysis. Volume 1: Power Series, Integration, Conformal Mapping, Location of Zeros. Wiley-Interscience, New York, NY, USA, 1974. ISBN 0-471-37244-7, 0-471-60841-6 (paperback). xv + 682 pp. LCCN QA331 .H453 1974.

Henrici:1974:CCA

[Hen74b] Peter Henrici. Computational complex analysis. In The influence of computing on mathematical research and education (Proceedings of the Symposium on Applied Mathematics, University of Montana, Missoula, Montana, 1973), volume 20, pages 79–86. American Mathematical Society, Providence, RI, USA, 1974.

Henrici:1975:EAK

[Hen75] Peter Henrici. Einige Anwendungen der Kreisscheibenarithmetik in der Kettenbruchtheorie. (German) [Some applications of interval arithmetic in continued fraction theory]. In Nickel [Nic75], pages 19–30. CODEN LNCSD9. ISBN 0-387-07170-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA297 .I541. DM30.00. URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0029.htm;   http://www.

springerlink.com/content/978-0-387-07170-1;   http://
www.springerlink.com/openurl.asp?genre=issue&issn=0302-
9743&volume=29;   https://link.springer.com/book/10.1007/

3-540-07170-9. German or English.

Henrici:1976:EAS

[Hen76] Peter Henrici. Einige Anwendungen der schnellen Fouriertransformation. (German) [Some applications of Fast Fourier Transform]. In Albrecht and Collatz [AC76], pages 111–124. Internat. Ser. Numer. Math., Vol. 32. ISBN 3-0348-5501-X. URL https://link.

springer.com/chapter/10.1007/978-3-0348-5501-3_8.

Henrici:1977:NTS

[Hen77a] P. Henrici. Note on a theorem of Saff and Varga. In Padé and rational approximation (Proceedings of the International Symposium, Universit of South Florida, Tampa, Florida, 1976), pages 157–161. Academic Press, New York, NY, USA, 1977.

Henrici:1977:ART

[Hen77b] Peter Henrici. Analytische Rechenverfahren für den Taschenrechner HP-25. (German) [Analytical Calculations for the HP-25 Calculator]. Oldenbourg, München, West Germany, 1977. ISBN 0-471-02938-6. 280 pp. LCCN ????

Henrici:1977:ACC

[Hen77c] Peter Henrici. Applied and Computational Complex Analysis. Volume 2: Special Functions, Integral Transforms, Asymptotics, Continued Fractions. Wiley, New York, NY, USA, 1977. ISBN 0-471-01525-3. ix + 662 pp.

Henrici:1977:EPD

[Hen77d] Peter Henrici. Error Propagation for Difference Methods. The SIAM series in applied mathematics. Robert E. Krieger Publishing Company, Huntington, NY, USA, 1977. ISBN 0-88275-448-3. vii + 73 pp. LCCN QA431 .H44; QA431 .H46 1977. Reprint of the 1963 edition. Companion to the author’s Discrete variable methods in ordinary differential equations.

Henrici:1978:BGB

[Hen78] P. Henrici. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 29:362–364, March 1978. CO-DEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

URL http://adsabs.harvard.edu/abs/1978ZaMP...29..362H.

Henrici:1979:NBF

[Hen79a] P. Henrici. Zur numerischen Berechnung der Fresnelschen Integrale. (German) [On numerical calculation of Fresnel integrals]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 30(2):209–219, 1979. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (electronic).

Henrici:1979:BFI

[Hen79b] Peter Henrici. Barycentric formulas for interpolating trigonometric polynomials and their conjugates. Numerische Mathematik, 33(2): 225–234, 1979. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Henrici:1979:BRA

[Hen79c] Peter Henrici. Book review: Accélération de la convergence en analyse numérique. Bulletin of the American Mathematical Society (new series), 1(4):639–641, 1979. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Henrici:1979:FFM

[Hen79d] Peter Henrici. Fast Fourier methods in computational complex analysis. SIAM Review, 21(4):481–527, October 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://www.jstor.org/stable/pdf/10.2307/2030105.pdf.

Henrici:1980:BRS

[Hen80a] Peter Henrici. Book review: The Symmetric Eigenvalue Problem, by Beresford N. Parlett, Prentice-Hall, Englewood Cliffs, NJ 1980, 348 pp. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 31(6):794, November 1980. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (electronic).

Henrici:1980:MPR

[Hen80b] Peter Henrici. A model for the propagation of rounding error in floating arithmetic. In Interval mathematics, 1980 (Freiburg, 1980), pages 49–73. Academic Press, New York, NY, USA, 1980.

ISBN 0-12-518850-1. URL http://www.sciencedirect.com/
science/article/pii/B9780125188500500093.

Henrici:1982:BGBa

[Hen82a] P. Henrici. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 33:149–150, January 1982. CO-DEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

URL http://adsabs.harvard.edu/abs/1982ZaMP...33..149H.

Henrici:1982:BGBc

[Hen82b] Peter Henrici. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 33(1):149–150, 1982. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Henrici:1982:DIP

[Hen82c] Peter Henrici. A definite integral: Problem 82-17. SIAM Review, 24(4):477, October 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https://www.jstor.org/

stable/2029543.

Henrici:1982:ENA

[Hen82d] Peter Henrici. Essentials of Numerical Analysis with Pocket Calculator Demonstrations. Wiley, New York, NY, USA, 1982. ISBN 0-471-05904-8. vi + 409 pp.

Henrici:1982:PHZ

[Hen82e] Peter Henrici. Peter Henrici: 91 Zitate aus den Jahren 1948–1982. (German) [Peter Henrici: 91 citations from the years 1948–1982]. Report 1224, ETH, Zürich, Switzerland, 1982. 11 pp.

Henrici:1982:PEE

[Hen82f] Peter Henrici. Pointwise error estimates for trigonometric interpolation and conjugation. Journal of Computational and Applied Mathematics, 8(2):131–132, 1982. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0771050X82900675.

Henrici:1983:GTO

[Hen83a] Peter Henrici. A general theory of osculation algorithms for conformal mapping. Linear Algebra and its Applications, 52/53: 361–382, 1983. CODEN LAAPAW. ISSN 0024-3795 (print),

1873-1856 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002437958380024X.

Henrici:1983:TCCa

[Hen83b] Peter Henrici. Topics in computational complex analysis. I. Methods of descent for polynomial equations. In Computational aspects of complex analysis (Braunlage, 1982), volume 102 of NATO Adv. Sci. Inst. Ser. C Math. Phys. Sci., pages 133–147. Reidel, Dordrecht, 1983.

Henrici:1983:TCCb

[Hen83c] Peter Henrici. Topics in computational complex analysis. II. New developments concerning the quotient-difference algorithm. In Computational aspects of complex analysis (Braunlage, 1982), volume 102 of NATO Adv. Sci. Inst. Ser. C Math. Phys. Sci., pages 149–168. Reidel, Dordrecht, 1983.

Henrici:1983:TCCc

[Hen83d] Peter Henrici. Topics in computational complex analysis. III. The asymptotic behavior of best approximations to analytic functions on the unit disk. In Computational aspects of complex analysis (Braunlage, 1982), volume 102 of NATO Adv. Sci. Inst. Ser. C Math. Phys. Sci., pages 169–191. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983.

Henrici:1983:TCCd

[Hen83e] Peter Henrici. Topics in computational complex analysis. IV. The Lagrange–Bürmann formula for systems of formal power series. In Computational aspects of complex analysis (Braunlage, 1982), volume 102 of NATO Adv. Sci. Inst. Ser. C Math. Phys. Sci., pages 193–215. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983.

Henrici:1984:LBF

[Hen84a] P. Henrici. Die Lagrange-Bürmannsche Formel bei formalen Potenzreihen. (German) [The Lagrange–Bürmann formula for formal power series]. Jahresbericht der Deutschen Mathematiker-Vereinigung, 86(4):115–134, 1984. CODEN JDMVA7. ISSN 0012-0456 (print), 1869-7135 (electronic).

Henrici:1984:RNI

[Hen84b] P. Henrici. Remark on numerical integration. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 35(5):712–714, 1984. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Henrici:1984:PEH

[Hen84c] Peter Henrici. Poisson’s equation in a hypercube: discrete Fourier methods, eigenfunction expansions, Padé approximation to eigenvalues. In Golub [Gol84], pages 371–411. ISBN 0-88385-126-1 (v. 1), 0-88385-100-8 (set). LCCN QA297 .S83 1984.

Henrici:1986:ACC

[Hen86a] Peter Henrici. Applied and Computational Complex Analysis. Volume 3: Discrete Fourier Analysis, Cauchy Integrals, Construction of Conformal Maps, Univalent Functions. Pure and Applied Mathematics (New York). Wiley, New York, NY, USA, 1986. ISBN 0-471-08703-3. xvi + 637 pp.

Henrici:1986:BRI

[Hen86b] Peter Henrici. Book review: Introduction to Applied Mathematics by Gilbert Strang. SIAM Review, 28(4):590–592, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL https://www.jstor.org/stable/2031128.

Henrici:1987:BPB

[Hen87a] Peter Henrici. De Branges’ proof of the Bieberbach conjecture: a view from computational analysis. In Sitzungsberichte der Berliner Mathematischen Gesellschaft, pages 105–121. Freie Universität Berlin, Berlin, West Germany, 1987.

Henrici:1987:TPT

[Hen87b] Peter Henrici. A triple product theorem for hypergeometric series. SIAM Journal on Mathematical Analysis, 18(6):1513–1518, 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Henrici:1988:ACC

[Hen88a] Peter Henrici. Applied and Computational Complex Analysis. Volume 1: Power Series, Integration, Conformal Mapping, Location of Zeros, volume 1 of Wiley Classics Library. Wiley, New York, NY, USA, 1988. ISBN 0-471-60841-6. xviii + 682 pp. Reprint of the 1974 original.

Henrici:1988:PTF

[Hen88b] Peter Henrici. Product theorems for formal hypergeometric series. In Visiting scholars’ lectures—1987 (Lubbock, TX), volume 15 of Texas Tech Univ. Math. Ser., pages 53–69. Texas Tech University, Lubbock, TX, USA, 1988.

Henrici:1989:SAP

[Hen89] Peter Henrici. Schnelle Algorithmen für periodische Funktionen. (German) [Fast algorithms for periodic functions]. Nova Acta Leopoldina (N.F.), 61(267):51–63, 1989. ISSN 0369-5034. Numerische Mathematik und ihre Anwendungen (Halle, 1982).

Henrici:1991:ACC

[Hen91] Peter Henrici. Applied and Computational Complex Analysis. Volume 2: Special Functions, Integral Transforms, Asymptotics, Continued Fractions, volume 2 of Wiley Classics Library. Wiley, New York, NY, USA, 1991. ISBN 0-471-54289-X. x + 662 pp. Reprint of the 1977 original.

Herriot:1966:RPP

[Her66] J. G. Herriot. Recent publications and presentations: Error Propagation for Difference Methods, by Peter Henrici. American Mathematical Monthly, 73(5):562, May 1966. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/2315512.

Henrici:1981:BGB

[HES^{+}81] P. Henrici, M. Eichler, E. Specker, R. Monnier, C. Blatter, R. Jost, and K. Hazeghi. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 32:357–360, May 1981. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL

http://adsabs.harvard.edu/abs/1981ZaMP...32..357H.

Henrici:1969:UCA

[HG69] Peter Henrici and Irene Gargantini. Uniformly convergent algorithms for the simultaneous approximation of all zeros of a polynomial. In Dejon and Henrici [DH69], pages 77–113. ISBN 0-471-20300-9. LCCN QA212 .C65.

Henrici:1975:CCC

[HH75] Peter Henrici and Christian Hoffmann. Chess champ’s chances: an exercise in asymptotics. SIAM Review, 17(3):559–564, July 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL https://www.jstor.org/stable/pdf/10.2307/
2028891.pdf.

Henrici:1981:PSS

[HH81] Peter Henrici and L. Van Hamme. Problems and solutions: Solutions of elementary problems: E2808. American Mathematical Monthly, 88(3):211, 1981. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

links.jstor.org/sici?sici=0002-9890(198103)88:3<211:E>
2.0.CO;2-X&origin=MSN;   https://www.jstor.org/stable/
2320479.

Henrici:1971:BGBb

[HHG^{+}71] P. Henrici, K. Hepp, M. Gutknecht, M. Sayir, K. Magnus, W. Epprecht, J. T. Marti, H. J. Lüthi, and F. Lenz. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 22: 996–1000, September 1971. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://adsabs.harvard.

edu/abs/1971ZaMP...22..996H.

Henrici:1977:KAI

[HJ80] Peter Henrici and Rita Jeltsch. Komplexe Analysis für Ingenieure. (German) [Complex Analysis for Engineers], volume 627 of Uni-Taschenbuecher. Birkhäuser Verlag, Basel, Switzerland, 1977– 1980. ???? pp. LCCN ???? URL https://link.springer.

com/book/10.1007/978-3-0348-5104-6.

Henrici:1982:BGBb

[HKC^{+}82] P. Henrici, A. Korányi, S. D. Chatterji, W. Traupel, F. Fankhauser, E. Engeler, and U. Kirchgraber. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 33:439– 442, May 1982. CODEN ZAMPDB. ISSN 0044-2275 (print),

1420-9039 (electronic). URL http://adsabs.harvard.edu/abs/
1982ZaMP...33..439H.

Henrici:1966:BGB

[HMC^{+}66] P. Henrici, M. Marchal, E. T. Cranch, H. Ziegler, W. Saxer, P. Grassmann, and C. Wehrli. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 17:654– 656, September 1966. CODEN ZAMPDB. ISSN 0044-2275 (print),

1420-9039 (electronic). URL http://adsabs.harvard.edu/abs/
1966ZaMP...17..654H.

Howlett:1966:BRE

[How66] J. Howlett. Book review: Elements of Numerical Analysis, by Peter Henrici. Mathematical Gazette, 50(373):357–358, October 1966. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.jstor.org/stable/3614752.

Henrici:1966:TEE

[HP66] Peter Henrici and Pia Pfluger. Truncation error estimates for Stieltjes fractions. Numerische Mathematik, 9(2):120–138, December 1966. CODEN NUMMA7. ISSN 0029-599x (print), 0945-

3245 (electronic). URL https://link.springer.com/article/
10.1007/BF02166031.

Henrici:1970:SFH

[HTW70] Peter Henrici, B. Andreas Troesch, and Luc Wuytack. Sloshing frequencies for a half-space with circular or strip-like aperture. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 21(3):285–318, May 1970. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (elec-

tronic). URL https://link.springer.com/article/10.1007/
BF01627938.

Henrici:1965:FZP

[HW65] P. Henrici and Bruce O. Watkins. Finding zeros of a polynomial by the Q–D algorithm. Communications of the ACM, 8:570– 574, 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrections [Tho66, Hen66b].

Henrici:1987:AML

[HW87] P. Henrici and G. R. Wilkens. An application of the multivariate Lagrange–Bürmann expansion in mathematical geodesy. Quarterly of Applied Mathematics, 45(1):165–172, 1987. CODEN QA-MAAY. ISSN 0033-569x (print), 1552-4485 (electronic).

Henrici:1968:ENC

[HZ68] Peter Henrici and Katharina Zimmermann. An estimate for the norms of certain cyclic Jacobi operators. Linear Algebra and its Applications, 1:489–501, 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0024379568900232.

Halg:1966:BGB

[HZP^{+}66] W. Hälg, H. Ziegler, P. Profos, H. R. Schwarz, A. Pfluger, T. Fischer, P. de Haller, and P. Henrici. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 17:548– 552, July 1966. CODEN ZAMPDB. ISSN 0044-2275 (print),

1420-9039 (electronic). URL http://adsabs.harvard.edu/abs/
1966ZaMP...17..548H.

I:1972:BRC

[I.72] E. I. Book review: Constructive Aspects of the Fundamental Theorem of Algebra, by B. Dejon and P. Henrici. Mathematics of Computation, 26(118):588–589, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.jstor.org/stable/2005189.

Kirchgraber:1981:BGB

[KHCW81a] U. Kirchgraber, P. Henrici, S. D. Chatterji, and V. Wüthrich. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 32:623–624, September 1981. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http:

//adsabs.harvard.edu/abs/1981ZaMP...32..623K.

Kirchgraber:1981:B

[KHCW81b] U. Kirchgraber, Peter Henrici, S. D. Chatterji, and V. Wüthrich. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 32(5):623–624, 1981. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Kunzi:1968:BGB

[KHE^{+}68] H. P. Künzi, P. J. Huber, W. Epprecht, I. Zschokke-Gränacher, K. Alder, P. Henrici, and E. Baldinger. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 19:154– 156, January 1968. CODEN ZAMPDB. ISSN 0044-2275 (print),

1420-9039 (electronic). URL http://adsabs.harvard.edu/abs/
1968ZaMP...19..154K.

Karlsson:1990:NHT

[KS90] Per W. Karlsson and H. M. Srivastava. A note on Henrici’s triple product theorem. Proceedings of the American Mathematical Society, 110(1):85, September 1990. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL https://www.jstor.

org/stable/2048245.

Leissa:1979:BGB

[LHL^{+}79] A. Leissa, F. Hampel, H. R. Lüthi, M. Dufour, and P. Henrici. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 30:871–872, September 1979. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http:

//adsabs.harvard.edu/abs/1979ZaMP...30..871L.

Loosli:1984:BGB

[LKM^{+}84] H. H. Loosli, M. Klemm, J. Marti, M. R. Trummer, P. Henrici, W. Lüthy, H. Giger, A. Friedli, J. J. Burckhardt, A. Friedli, H. Brunner, and H. R. Schwarz. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 35:920– 924, November 1984. CODEN ZAMPDB. ISSN 0044-2275 (print),

1420-9039 (electronic). URL http://adsabs.harvard.edu/abs/
1984ZaMP...35..920L.

Macon:1965:ENA

[Mac65a] Nathaniel Macon. Book review: Elements of Numerical Analysis, by Peter Henrici. American Mathematical Monthly, 72(10): 1151–1152, December 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/2315994.

Macon:1965:RPP

[Mac65b] Nathaniel Macon. Recent publications and presentations: Elements of Numerical Analysis, by Peter Henrici. American Mathematical Monthly, 72(10):1151–1152, December 1965. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

MacVean:1968:BGB

[MHZ^{+}68] D. MacVean, P. Henrici, H. Ziegler, R. Zuber, M. Sayir, P. Schmid, M. Gaillard, and P. Huguenin. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 19:378– 380, March 1968. CODEN ZAMPDB. ISSN 0044-2275 (print),

1420-9039 (electronic). URL http://adsabs.harvard.edu/abs/
1968ZaMP...19..378M.

Moore:1962:BRD

[Moo62] Ramon E. Moore. Book review: Discrete Variable Methods in Ordinary Differential Equations by Peter Henrici. Science, 136(3511):143–144, April 13, 1962. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

science.sciencemag.org/content/136/3511/143.2;   http:
//www.jstor.org/stable/1708446.

Morris:1971:CAN

[Mor71] John Ll. Morris, editor. Conference on Applications of Numerical Analysis: Held in Dundee/Scotland, March 23–26, 1971, volume 228 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L28 no. 228. URL http://link.springer.com/book/10.1007/
BFb0069442;   http://www.springerlink.com/content/978-
3-540-36976-9.

Nickel:1967:ERP

[Nic67] Karl L. E. Nickel. Extension of a recent paper by Fox, Henrici and Moler on eigenvalues of elliptic operators. SIAM Journal on Numerical Analysis, 4(4):483–488, December 1967. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL

https://www.jstor.org/stable/2949713.

Nickel:1975:IMP

[Nic75] Karl Nickel, editor. Interval mathematics: proceedings of the international symposium, Karlsruhe, West Germany, May 20–24, 1975, volume 29 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNCSD9. ISBN 0-387-07170-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA297 .I541. DM30.00. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t0029.htm;   http://www.springerlink.com/content/978-
0-387-07170-1;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=29;   https://link.
springer.com/book/10.1007/3-540-07170-9. German or En-

glish.

O:1964:BRBa

[O.64] I. O. Book review: Error Propagation for Difference Methods, by Peter Henrici. Journal of the American Statistical Association, 59(305):314, March 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2282896;   https://www.jstor.org/
stable/2282897.

OConnor:2008:PKH

[OR08] J. J. O’Connor and E. F. Robertson. Peter K. Henrici. Web site,

July 2008. URL http://www-history.mcs.st-andrews.ac.uk/
Biographies/Henrici_Peter.html.

Oyibo:1986:BRA

[Oyi86] G. A. Oyibo. Book review: Applied and Computational Complex Analysis, Vol. 3 , by Peter Henrici. American Institute of Aeronautics and Astronautics Journal, 24:2055–2056, December 1986. CODEN AIAJAH. ISSN 0001-1452. URL http://adsabs.

harvard.edu/abs/1986AIAAJ..24.2055O.

Pfluger:1966:TEE

[PH66] A. Pfluger and P. Henrici. Truncation error estimates for Stieltjes fractions. In Contemporary Problems in Theory Anal. Functions (International Conference, Erevan, 1965) (Russian), pages 257– 262. Izdatel’stvo “Nauka”, Moscow, USSR, 1966.

Rhanizar:2022:MVH

[Rha22] Bouchta Rhanizar. On a modified version of the Henrici’s transformation into optimization. Numerical Algorithms, 90(2):507–522, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01196-5.

S:1978:BRM

[S.78] H. J. S. Book review: Martin Gutknecht, Peter Henrici, Peter Lauchli, and Hans-Rudolf Schwarz, Heinz Rutishauser: Vorlesungen Uber Numerische Mathematik Vols. 1 and 2. Mathematics of Computation, 32(142):651–652, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.jstor.org/stable/pdf/2006176.pdf.

S:1987:BRA

[S.87] F. S. Book review: Applied and Computational Complex Analysis, Vol. 3: Discrete Fourier Analysis–Cauchy Integrals–Construction of Conformal Maps–Univalent Functions, by Peter Henrici. Mathematics of Computation, 48(178):844–845, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

https://www.jstor.org/stable/2007855.

Sadok:1990:AHT

[Sad90] Hassane Sadok. About Henrici’s transformation for accelerating vector sequences. Journal of Computational and Applied Mathematics, 29(1):101–110, January 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037704279090199A.

Shenton:1979:PSE

[SBL^{+}79] L. R. Shenton, Frank Bowman, H. K. Lam, Harry D. Ruderman, Wells Johnson, F. S. Cater, Solomon W. Golomb, and P. Henrici. Problems and solutions: Elementary problems: E2803–E2808. American Mathematical Monthly, 86(10):864–865, 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Seewald:1985:BGB

[SBM^{+}85] W. Seewald, T. Binkert, Q. Munir, U. Kirchgraber, W. Filter, P. Henrici, Z. Kunszt, C. Blatter, R. Sperb, E. Straube, and T. M. Liebling. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 36:332–336, March 1985. CO-DEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

URL http://adsabs.harvard.edu/abs/1985ZaMP...36..332S.

Schechter:1966:BRD

[Sch66a] Samuel Schechter. Book review: Discrete Variable Methods in Ordinary Differential Equations, by Peter Henrici. Mathematics of Computation, 20(93):183–184, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.jstor.org/stable/pdf/2004310.pdf.

Schmidt:1966:BPHa

[Sch66b] J. W. Schmidt. Buchbesprechung: P. Henrici, Discrete Variable Methods in Ordinary Differential Equations xi + 407 S. m. Abb. u.

Tab. New York/London 1962. John Wiley & Sons. Zeitschrift für Angewandte Mathematik und Mechanik, 46(5):327–328, 1966. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Schmidt:1966:BPHb

[Sch66c] J. W. Schmidt. Buchbesprechung: P. Henrici, Error Propagation for Difference. (The SIAM Series in Applied Mathematics) 72 S. u. 8. Fig. New York/London 1963. John Wiley and Sons, Inc. Zeitschrift für Angewandte Mathematik und Mechanik, 46(5):328, 1966. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Straumann:1980:BGB

[SMO^{+}80] N. Straumann, G. Mislin, J. L. Olsen, P. Henrici, K. Amsler, K. Marti, A. Piaget, M. von Allmen, A. Moser, K. Meier, P. Henrici, A. Furrer, U. Stammbach, H. Ziegler, J. Hersch, G. Schweitzer, H.-J. Lüthi, J. J. Burckhardt, E. Engeler, M. A. Knus, P. Läuchli, R. Schilling, and C. Blatter. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 31:793–802, November 1980. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://adsabs.harvard.

edu/abs/1980ZaMP...31..793S.

Sterbenz:1974:FPC

[Ste74] Pat H. Sterbenz. Floating Point Computation. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1974. ISBN 0-13-322495-3. xiv + 316 pp. LCCN QA76.8.I12 S771 1974.

Stutz:1964:BGB

[SZR^{+}64] T. Stutz, H. Ziegler, E. Roth, E. Baldinger, B. Thürlimann, K. Alder, H. M. Fischer, G. Busch, W. Saxer, P. Läuchli, P. de Haller, and P. Henrici. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 15:662–666, November 1964. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039

(electronic). URL http://adsabs.harvard.edu/abs/1964ZaMP.
..15..662S.

Tan:1995:HTA

[TA95] Roger C. E. Tan and Alan L. Andrew. Henrici’s transformation and its application to the computation of derivatives of eigensystems. Journal of Computational and Applied Mathematics, 59(1):

1–8, April 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037704279400013Q.

Traupel:1968:BGB

[TBH^{+}68] W. Traupel, E. Baldinger, P. Henrici, M. Sayir, H. Thomann, and H. R. Schwarz. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 19:541–544, May 1968. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL

http://adsabs.harvard.edu/abs/1968ZaMP...19..541T.

Troyon:1986:BGB

[TBS^{+}86] F. Troyon, C. Bandle, H. R. Schwarz, G. Seeber, C. Bandle, J. O. Stenflo, H.-J. Lüthi, and P. Henrici. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 37:451– 454, May 1986. CODEN ZAMPDB. ISSN 0044-2275 (print),

1420-9039 (electronic). URL http://adsabs.harvard.edu/abs/
1986ZaMP...37..451T.

Thomann:1979:BGB

[TGH^{+}79] H. Thomann, A. Gyr, P. Henrici, M. Monkewitz, C. Blatter, G. Schüler, P. Profos, E. Engeler, P. Läuchli, J. Kemper, A. Friedli, G. Gyarmathy, N. Rott, and T. M. Liebling. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 30: 1030–1036, November 1979. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://adsabs.harvard.

edu/abs/1979ZaMP...30.1030T.

Thacher:1977:ACC

[Tha77] Henry C. Thacher. Book review: Applied and Computational Complex Analysis, by Peter Henrici. Mathematics of Computation, 31(137):325–326, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.jstor.

org/stable/pdf/2005805.pdf.

Thacher:1981:CAH

[Tha81] Henry C. Thacher. Book review: Computational Analysis with the HP-25 Pocket Calculator, by Peter Henrici. Mathematics of Computation, 37(156):595–596, October 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://

www.jstor.org/stable/2007449.

Thomas:1966:LEC

[Tho66] Richard F. Thomas, Jr. Letter to the Editor: Corrections to numerical data on Q–D algorithm. Communications of the ACM, 9(5):322–323, May 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [HW65, Hen66b].

Uhlig:1966:BHP

[Uhl66] J. Uhlig. Buchbesprechung: Henrici, P. Elementes of Numerical Analysis, xv + 328 S. m. Fig. New York/London/Sydney 1964. John Wiley & Sohn. Preis geb. 60 s. net. Zeitschrift für Angewandte Mathematik und Mechanik, 46(5):329, 1966. CODEN ZA-MMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

VanTuyl:1983:BRB

[VJ83] Andrew H. Van Tuyl and Harry L. Johnson. Book review: A Definite Integral (Peter Henrici). SIAM Review, 25(4):572–573, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vogel:1979:PSE

[VPC^{+}79] John W. Vogel, Robert Patenaude, E. D. Camier, Robert A. Leslie, Doug Wiedemann, and P. Henrici. Problems and solutions: Elementary problems: E2791–E2796. American Mathematical Monthly, 86(8):702–703, 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Willoughby:1964:BRB

[Wil64] Ralph A. Willoughby. Book review: Error Propagation for Difference Methods (Peter Henrici). SIAM Review, 6(3):316–317, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL https://www.jstor.org/stable/2027606.

Willoughby:1965:BRB

[Wil65] Ralph A. Willoughby. Book review: Elements of Numerical Analysis (Peter Henrici). SIAM Review, 7(4):571–572, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL https://www.jstor.org/stable/2027792.

Waldvogel:1979:ES

[WKSH79] J. Waldvogel, U. Kirchgraber, H. R. Schwarz, and P. Henrici. Eduard Stiefel (1909–1978). Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 30(2):

133–142 + 1, 1979. CODEN ZAMPA8. ISSN 0044-2275 (print), 1420-9039 (electronic). With a bibliography and comments on the bibliography by Stiefel.

Ziegler:1966:BGB

[ZFL^{+}66] H. Ziegler, M. Fierz, A. Linder, W. Saxer, and P. Henrici. Buchbesprechungen. (German) [Book reviews]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 17:487–488, May 1966. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic). URL http://adsabs.

harvard.edu/abs/1966ZaMP...17..487Z.